Hint: Recall that speed is distance/time 2. [2pt] A car travels a certain distance along a straight road (in the +x direction someone answered this previously but it was incorrect they had TTFTEL
Answer #1

EGGTGL

A) After roughly extrapolating at time=60s and time=80s, the speed comes out to be the same. Also since the slope is linear you can say that acceleration is equal to 0.

B) Speed continuously decreases as you can see. S1= 300/20, S2=300/30. So, speed is still greater than zero.

C) The graph seems to be increasing again at t=37s. So here acceleration is greater than 0.

D) After t= 60s, the graph is linear till t=80s, meaning the speed is the same and does not increase further.

E) At t=0, acceleration is zero and at t=37s acceleration is slightly greater than 0.

F) From the slope of the graph, it is clear that speed at 70sec is less than the speed at 4sec because the slope is higher in the range t=0 to t=20s meaning higher speed than the speed in the range t=60s to t=80s.